# PaperQA: Retrieval-Augmented Generative Agent for Scientific Research

### 08 Dec 2023 v1

@@ 1 Introduction
The rate of papers published yearly grows at an exponential rate, with over 5 million academic articles published in 2022 (Curcic, 2023), and over 200 million articles in total (Fire and Guestrin, 2019). The difficulty of navigating this extensive literature means significant scientific findings have gone unnoticed for extended periods (Fortunato et al., 2018, Stent, 1972, Garfield, 1980). Work in the last 10 years has sought to make the space of literature more manageable for scientists, with the introduction of keyword search systems (Google Scholar, , Kinney et al., 2023), vector similarity embeddings (Cohan et al., 2020, Beltagy et al., 2019, Bojanowski et al., 2017, Singh et al., 2022) and recommender systems (Resnick and Varian, 1997, Adomavicius and Tuzhilin, 2005). The process of scientific discovery from literature is still, however, highly manual.
The use of Large Language Models (LLMs) to answer scientific questions is increasingly seen in academia and research-heavy professions such as medicine (Singhal et al., 2023, Williams et al., 2023). While LLMs can produce answers faster and encompass a broader, deeper scope than manual searching, there is a high risk of hallucination in responses, which can lead to potentially dangerous outcomes (Hiesinger et al., 2023, Meskó and Topol, 2023). Incorrect information can be more damaging than no information at all, as the time to verify veracity can take just as long as retrieving it from research papers in the first place (Williamson, 2016). Reliance on pre-trained LLMs also prevents the discovery of new information published after a training cutoff date. Given the rapidly moving pace of science, this can lead to misconceptions persisting.
Retrieval-Augmented Generation (RAG) models are a potential solution to these limitations (Lewis et al., 2020b). RAG models retrieve text from a corpus, using methods such as vector embedding search or keyword search, and add the retrieved passage to the context window of the LLM. RAG usage can reduce hallucinations in conversations and improve LLM performance on QA tasks (Hiesinger et al., 2023, Shuster et al., 2021). Nevertheless, standard RAG models follow a fixed, linear flow, which can be restrictive for addressing the diverse range of questions scientists encounter.
@@
We evaluate PaperQA on several standard multiple-choice datasets for evaluating LLMs, which assess the model’s ability to answer questions with existing knowledge, but do not test the agents’ ability to retrieve information. We modified PubMedQA Jin et al. (2019) to remove the provided context (so it is closed-book) and found PaperQA beats GPT-4 by 30 points (57.9% to 86.3%). PubMedQA is only built on abstracts though, and so we construct a more difficult dataset that requires synthesizing information from one or multiple full-text research papers. We thus introduce a new dataset, LitQA, composed from recent literature, in order to test PaperQA’s ability to retrieve information outside of the underlying LLM’s pre-training data. PaperQA outperforms all models tested and commercial tools, and is comparable to human experts on LitQA on performance and time, but is significantly cheaper in terms of costs. Furthermore, PaperQA is competitive to state-of-the-art commercial tools for scientific question answering. Finally, we find that PaperQA exhibits a better knowledge boundary than competing tools, answering questions incorrectly at a lower rate, and instead, answering that it is unsure.
@@ 2 Related Works
Large Language Models (LLMs) have seen a surge in popularity and accessibility, leading to impressive applications across various domains (Vaswani et al., 2017, Brown et al., 2020, Bommasani et al., 2021, Chowdhery et al., 2022, Driess et al., 2023, OpenAI, 2023). LLMs have been used effectively for text-based scientific tasks, such as extracting chemical reaction procedures (Bai et al., 2022) and entity extraction from biological documents (Tamari et al., 2021). LLMs trained on biomedical (Gu et al., 2021, Lewis et al., 2020a, Luo et al., 2022, Lee et al., 2020, Shin et al., 2020) or more generalized scientific literature (Beltagy et al., 2019, Taylor et al., 2022) have demonstrated impressive performance on current scientific QA benchmarks.
State-of-the-art pre-trained LLMs exhibit proficiency in complex tasks such as searching for chemical compounds (OpenAI, 2023) and designing new catalysts (Ramos et al., 2023). Still, LLMs frequently fall short in many scientific domains due to outdated knowledge (Fan et al., 2022) and reasoning problems (Shinn et al., 2023, Wang et al., 2022). Prompt engineering techniques can enhance LLMs’ reasoning abilities (Wei et al., 2022, Yao et al., 2023), though some scientific tasks requiring real-time calculations and up-to-date information remain difficult.
@@
Another technique is to integrate external tools into LLMs, creating agent systems, such as MRKL Karpas et al. (2022). These LLM-agent systems leverage the reasoning power of the base-LLM and the use of pre-defined external tools to act in order to complete a given prompt Schick et al. (2023), Shen et al. (2023). In such a system, the agent iteratively decides on the best tools to use in order to address the given task, correcting previous behavior when necessary, until the task is complete. This setup can also integrate reasoning steps, as with the ReAct framework and its derivatives (Yao et al., 2022, Yang et al., 2023b), or may incorporate multi-agent systems (Talebirad and Nadiri, 2023, Wang et al., 2023, Significant-gravitas, 2023). Our method extends previous work by integrating an agent-based RAG framework in order to correctly and dependably answer scientific questions.
@@
Assessing the scientific capabilities of LLMs often relies on QA benchmarks, such as general science benchmarks (Kembhavi et al., 2017, Hendrycks et al., 2020), or those specializing in medicine (Jin et al., 2019), biomedical science (Sung et al., 2021) or chemistry (Guo et al., 2023, Wu et al., 2017). In contrast, open-ended tasks, such as conducting chemical synthesis planning (Bran et al., 2023) or offering healthcare support (Dash et al., 2023), necessitate manual evaluation to effectively measure an LLM’s capabilities. We discuss the limitation of these datasets and introduce a new QA dataset for retrieval-based science.
@@
LLMs integrated with retrieval systems are broadly referred to as Retrieval-Augmented Generation (RAG) models (Lewis et al., 2020b), with the key components being a database of documents, a query system for retrieving documents, and a pipeline to add retrieved documents to a model’s context. RAG models have proved effective at biomedical and clinical QA tasks (Soong et al., 2023, Hiesinger et al., 2023). While RAG systems are typically pipelines with a fixed number and order of steps, they can be composed of multi-hop searches (Khattab et al., 2022), or use Agents (Significant-gravitas, 2023), where an LLM is used to decide when to use retrieval to augment an answer (Schick et al., 2023). RAG performance can be further improved with a priori prompting, where an LLM is prompted to consider using latent information instead of searching, or a posteriori prompting, where an LLM is asked to consider the accuracy of a retrieved result and possibly provide an alternative answer (Ren et al., 2023). Active RAG (Jiang et al., 2023) regenerates sentences using retrieval if they contain low-probability tokens.
@@
RAG models connected to databases retrieve documents using fixed representations, such as Bag-of-Words or BM25 (Izacard and Grave, 2020, Doostmohammadi et al., 2023)), pre-trained embeddings (Devlin et al., 2018, Izacard and Grave, 2020), or trainable encoders (Lewis et al., 2020b, Shi et al., 2023), which are trained offline (Shi et al., 2023) or online (Guu et al., 2020). For models using open-ended sources such as the internet, phrase or keyword searches are used for retrieval (Menick et al., 2022, Significant-gravitas, 2023). Retrieved documents are typically passed to the model as input tokens, but some models separately process retrieved information with cross-attention (Borgeaud et al., 2022, Doostmohammadi et al., 2023). Learning to rank and filter the retrieved documents (Lyu et al., 2023) further improves the performance of RAG models. In our work, we evaluate Retrieval-Augmented Models that use keyword search systems and vector embedding retrieval. In particular, we focus on benchmarking against humans and uncover under what conditions LLMs can match human performance in information retrieval.
@@ 3.1 PaperQA
The PaperQA system is an agent shown in Figure 1. The fundamental operations of PaperQA are to find relevant papers from online databases, such as Arxiv [^1] https://arxiv.org/ and Pubmed [^2] https://pubmed.ncbi.nlm.nih.gov/; gather text from these papers; and synthesize information into a final answer. PaperQA’s tools are composed of both external APIs and LLMs, described in detail below. Compared to a standard RAG, we make four key innovations. First, we decompose parts of a RAG and provide them as tools to an agent. This allows for some operations, such as search, to be performed multiple times, with different keywords, if the evidence collected is insufficient. Second, we use a map summarization step to gather evidence from multiple sources followed by a reduce step to answer. The map-reduce step increases the amount of sources that can be considered, and provides a scratchpad (Wei et al., 2022), where the LLM can give intermediate evidence before formulating the final answer. Third, we utilise the LLM’s ability to reason over text and provide numerical relevance scores for each chunk of text to the query question. In addition to the commonly used vector-embedding distances, we use these LLM-generated relevancy scores, adding another layer of retrieval. Finally, we make use of a priori and a posteriori prompting, tapping into the latent knowledge in LLMs.
@@ 3.2 Tools
search The agent gives this tool keywords and (optionally) a year range. This queries a scientific literature search engine. Retrieved papers are added to a local bibliography for use by gather_evidence. They are added by creating overlapping 4,000 character chunks which are embedded with the text-embedding-ada-002 model (Neelakantan et al., 2022) and inserted into a vector database (Johnson et al., 2019). There is a failure rate associated with the performance of search engines, accessing papers, and parsing of PDFs. We explore this in Appendix C. This tool is always executed once with the full-text question before initialising the agent.
gather_evidence This tool is given a question as input by the agent. This question is vector-embedded using the OpenAI text-embedding-ada-002 (Neelakantan et al., 2022), and relevant chunks based on vector search are returned from the vector database created in the search tool. Chunks are retrieved with maximal marginal relevance search (Carbonell and Goldstein, 1998) to increase the diversity of returned texts. Each retrieved chunk is fed to a summary LLM with the following prompt:
@@
answer_question We first use the ask LLM to provide any useful information from the pre-trained LLM that might help with answering the original question (details are given in Appendix A.4), similar to the to the priori judgement explored in (Ren et al., 2023). The output of the ask LLM is added to the chunks from the context library, and the final prompt to the answer LLM is as follows:
Write an answer (answer_length) for the question below based on the provided context. If the context provides insufficient information, reply "I cannot answer". For each part of your answer, indicate which sources most support it via valid citation markers at the end of sentences, like (Example2012). Answer in an unbiased, comprehensive, and scholarly tone. If the question is subjective, provide an opinionated answer in the concluding 1-2 sentences.
@@ 5.1 Experimental Details
All of the experiments were implemented within LangChain’s agent framework (Chase, 2022). As described above, we use four different LLM instances with the following settings unless stated otherwise: agent LLM: GPT-4 OpenAIFunctions Agent with temperature $\tau_{\mathrm{agent}}=0.5$; summary LLM: GPT-3.5 always with temperature $\tau_{\mathrm{sum}}=0.2$; answer LLM: GPT-4 with temperature $\tau_{\mathrm{ans}}=0.5$; ask LLM: GPT-4 with temperature $\tau_{\mathrm{ask}}=0.5$. For the search engine, we use Google Scholar, where we collect the top-5 papers that are accessible. The implementation details are explained in an experiment in Appendix C. We access papers through publicly available APIs, such as Arxiv, PMC, OpenAccess, PubMed, and our own local database of papers. Lastly, we set the number of sources to consider at each round of gather_evidence to 20, and the count of evidence context to include in the final answer within answer_question to 8.
@@ 6 Limitations
While recent work has been conducted on prompt optimization (Deng et al., 2022, Yang et al., 2023a), the complex setting of multiple agents with individual prompts is unsolved. Specifically, the task becomes a non-trivial, bi-level optimization problem. Consequently, discerning the impact of manual prompt adjustments becomes difficult. Thus, it is unlikely our prompts are optimal and it is difficult to assess which pieces of the prompts are necessary.
@@ Acknowledgments
Dual-use (applying technology for both beneficial and detrimental purposes) is an ongoing concern as more powerful models are developed Urbina et al. (2022). We performed rigorous red-teaming questions and did not observe risks for harm that are significantly elevated compared to direct use of the language models used for summarization. Similarly, the LitQA dataset and responses pose no risk for harm.
@@ Abstract:
Provide num_keywords unique keyword searches (one search per line) and year ranges that will find papers to help answer the question. Do not use boolean operators. "Make sure not to repeat searches without changing the keywords or year ranges. Make some searches broad and some narrow. The current year is get_year().
@@
We thus evaluate recall, i.e. finding the original paper. Figure 2 shows the cumulative recall curve, where Google Scholar and Semantic Scholar show outstanding ability to retrieve the original paper. Although CORE API (Knoth et al., 2023) does not perform on par with the other two, we include it here as their main contribution to the field of scientific literature is their standardisation of open-access article repositories.
@@ Appendix G Impact of Parametric Knowledge
Write an answer (about 200 words, but can be longer if necessary) for the question below based on the provided context. If the context provides insufficient information and the question cannot be directly answered, reply "I cannot answer". For each part of your answer, indicate which sources most support it via valid citation markers at the end of sentences, like (Example2012).
@@ G.1 Contradicting Information
In the first example, we ask Are COVID-19 vaccines effective?. The goal of this experiment is to evaluate the inclusion of background information that contradicts gathered context. Thus, while the provided context supports the efficacy of covid vaccines, we provide the model with the following background information: ”COVID-19 vaccines are known to be ineffective.”
With the contradicting background information, PaperQA responds that it cannot answer. However, if the extra, contradicting information is excluded, the model provides an answer:
@@ G.2 Absence of Key Information
As a second example, we ask PaperQA What temperature does water turn to plasma? and provide it with context that is relevant but insufficient to answer the question. Then, we provide the model with the following background information:
”Water turns into plasma under extreme conditions of temperature and pressure. Plasma is the fourth state of matter, beyond solid, liquid, and gas. It’s a super-heated state where atoms are ionized, meaning electrons are stripped from atoms. This typically occurs at temperatures exceeding 20,000 Kelvin.”

### 14 Dec 2023 v2

@@ 1 Introduction
The rate of papers published yearly grows at an exponential rate, with over 5 million academic articles published in 2022 [1], and over 200 million articles in total [2]. The difficulty of navigating this extensive literature means significant scientific findings have gone unnoticed for extended periods [3, 4, 5]. Work in the last 10 years has sought to make the space of literature more manageable for scientists, with the introduction of keyword search systems [6, 7], vector similarity embeddings [8, 9, 10, 11] and recommender systems [12, 13]. The process of scientific discovery from literature is still, however, highly manual.
The use of Large Language Models (LLMs) to answer scientific questions is increasingly seen in academia and research-heavy professions such as medicine [14, 15]. While LLMs can produce answers faster and encompass a broader, deeper scope than manual searching, there is a high risk of hallucination in responses, which can lead to potentially dangerous outcomes [16, 17]. Incorrect information can be more damaging than no information at all, as the time to verify veracity can take just as long as retrieving it from research papers in the first place [18]. Reliance on pre-trained LLMs also prevents the discovery of new information published after a training cutoff date. Given the rapidly moving pace of science, this can lead to misconceptions persisting.
Retrieval-Augmented Generation (RAG) models are a potential solution to these limitations [19]. RAG models retrieve text from a corpus, using methods such as vector embedding search or keyword search, and add the retrieved passage to the context window of the LLM. RAG usage can reduce hallucinations in conversations and improve LLM performance on QA tasks [16, 20]. Nevertheless, standard RAG models follow a fixed, linear flow, which can be restrictive for addressing the diverse range of questions scientists encounter.
@@
We evaluate PaperQA on several standard multiple-choice datasets for evaluating LLMs, which assess the model’s ability to answer questions with existing knowledge, but do not test the agents’ ability to retrieve information. We modified PubMedQA [21] to remove the provided context (so it is closed-book) and found PaperQA beats GPT-4 by 30 points (57.9% to 86.3%). PubMedQA is only built on abstracts though, and so we construct a more difficult dataset that requires synthesizing information from one or multiple full-text research papers. We thus introduce a new dataset, LitQA, composed from recent literature, in order to test PaperQA’s ability to retrieve information outside of the underlying LLM’s pre-training data. PaperQA outperforms all models tested and commercial tools, and is comparable to human experts on LitQA on performance and time, but is significantly cheaper in terms of costs. Furthermore, PaperQA is competitive to state-of-the-art commercial tools for scientific question answering. Finally, we find that PaperQA exhibits a better knowledge boundary than competing tools, answering questions incorrectly at a lower rate, and instead, answering that it is unsure.
@@ 2 Related Works
Large Language Models (LLMs) have seen a surge in popularity and accessibility, leading to impressive applications across various domains [22, 23, 24, 25, 26, 27]. LLMs have been used effectively for text-based scientific tasks, such as extracting chemical reaction procedures [28] and entity extraction from biological documents [29]. LLMs trained on biomedical [30, 31, 32, 33, 34] or more generalized scientific literature [9, 35] have demonstrated impressive performance on current scientific QA benchmarks.
State-of-the-art pre-trained LLMs exhibit proficiency in complex tasks such as searching for chemical compounds [27] and designing new catalysts [36]. Still, LLMs frequently fall short in many scientific domains due to outdated knowledge [37] and reasoning problems [38, 39]. Prompt engineering techniques can enhance LLMs’ reasoning abilities [40, 41], though some scientific tasks requiring real-time calculations and up-to-date information remain difficult.
@@
Another technique is to integrate external tools into LLMs, creating agent systems, such as MRKL [42]. These LLM-agent systems leverage the reasoning power of the base-LLM and the use of pre-defined external tools to act in order to complete a given prompt [43, 44]. In such a system, the agent iteratively decides on the best tools to use in order to address the given task, correcting previous behavior when necessary, until the task is complete. This setup can also integrate reasoning steps, as with the ReAct framework and its derivatives [45, 46], or may incorporate multi-agent systems [47, 48, 49]. Our method extends previous work by integrating an agent-based RAG framework in order to correctly and dependably answer scientific questions.
@@
Assessing the scientific capabilities of LLMs often relies on QA benchmarks, such as general science benchmarks [50, 51], or those specializing in medicine [21], biomedical science [52] or chemistry [53, 54]. In contrast, open-ended tasks, such as conducting chemical synthesis planning [55] or offering healthcare support [56], necessitate manual evaluation to effectively measure an LLM’s capabilities. We discuss the limitation of these datasets and introduce a new QA dataset for retrieval-based science.
@@
LLMs integrated with retrieval systems are broadly referred to as Retrieval-Augmented Generation (RAG) models [19], with the key components being a database of documents, a query system for retrieving documents, and a pipeline to add retrieved documents to a model’s context. RAG models have proved effective at biomedical and clinical QA tasks [57, 16]. While RAG systems are typically pipelines with a fixed number and order of steps, they can be composed of multi-hop searches [58], or use Agents [49], where an LLM is used to decide when to use retrieval to augment an answer [43]. RAG performance can be further improved with a priori prompting, where an LLM is prompted to consider using latent information instead of searching, or a posteriori prompting, where an LLM is asked to consider the accuracy of a retrieved result and possibly provide an alternative answer [59]. Active RAG [60] regenerates sentences using retrieval if they contain low-probability tokens.
@@
RAG models connected to databases retrieve documents using fixed representations, such as Bag-of-Words or BM25 [61, 62], pre-trained embeddings [63, 61], or trainable encoders [19, 64], which are trained offline [64] or online [65]. For models using open-ended sources such as the internet, phrase or keyword searches are used for retrieval [66, 49]. Retrieved documents are typically passed to the model as input tokens, but some models separately process retrieved information with cross-attention [67, 62]. Learning to rank and filter the retrieved documents [68] further improves the performance of RAG models. In our work, we evaluate Retrieval-Augmented Models that use keyword search systems and vector embedding retrieval. In particular, we focus on benchmarking against humans and uncover under what conditions LLMs can match human performance in information retrieval.
@@ 3.1 PaperQA
The PaperQA system is an agent shown in Figure 1. The fundamental operations of PaperQA are to find relevant papers from online databases, such as Arxiv [^1] https://arxiv.org/ and Pubmed [^2] https://pubmed.ncbi.nlm.nih.gov/; gather text from these papers; and synthesize information into a final answer. PaperQA’s tools are composed of both external APIs and LLMs, described in detail below. Compared to a standard RAG, we make four key innovations. First, we decompose parts of a RAG and provide them as tools to an agent. This allows for some operations, such as search, to be performed multiple times, with different keywords, if the evidence collected is insufficient. Second, we use a map summarization step to gather evidence from multiple sources followed by a reduce step to answer. The map-reduce step increases the amount of sources that can be considered, and provides a scratchpad [40], where the LLM can give intermediate evidence before formulating the final answer. Third, we utilise the LLM’s ability to reason over text and provide numerical relevance scores for each chunk of text to the query question. In addition to the commonly used vector-embedding distances, we use these LLM-generated relevancy scores, adding another layer of retrieval. Finally, we make use of a priori and a posteriori prompting, tapping into the latent knowledge in LLMs.
@@ 3.2 Tools
search The agent gives this tool keywords and (optionally) a year range. This queries a scientific literature search engine. Retrieved papers are added to a local bibliography for use by gather_evidence. They are added by creating overlapping 4,000 character chunks which are embedded with the text-embedding-ada-002 model [69] and inserted into a vector database [70]. There is a failure rate associated with the performance of search engines, accessing papers, and parsing of PDFs. We explore this in Appendix C. This tool is always executed once with the full-text question before initialising the agent.
gather_evidence This tool is given a question as input by the agent. This question is vector-embedded using the OpenAI text-embedding-ada-002 [69], and relevant chunks based on vector search are returned from the vector database created in the search tool. Chunks are retrieved with maximal marginal relevance search [71] to increase the diversity of returned texts. Each retrieved chunk is fed to a summary LLM with the following prompt:
@@
answer_question We first use the ask LLM to provide any useful information from the pre-trained LLM that might help with answering the original question (details are given in Appendix A.4), similar to the to the priori judgement explored in [59]. The output of the ask LLM is added to the chunks from the context library, and the final prompt to the answer LLM is as follows:
Write an answer (answer_length) for the question below based on the provided context. If the context provides insufficient information, reply ‘‘I cannot answer’’. For each part of your answer, indicate which sources most support it via valid citation markers at the end of sentences, like (Example2012). Answer in an unbiased, comprehensive, and scholarly tone. If the question is subjective, provide an opinionated answer in the concluding 1-2 sentences.
@@ 5.1 Experimental Details
All of the experiments were implemented within LangChain’s agent framework [72]. As described above, we use four different LLM instances with the following settings unless stated otherwise: agent LLM: GPT-4 OpenAIFunctions Agent with temperature $\tau_{\mathrm{agent}}=0.5$; summary LLM: GPT-3.5 always with temperature $\tau_{\mathrm{sum}}=0.2$; answer LLM: GPT-4 with temperature $\tau_{\mathrm{ans}}=0.5$; ask LLM: GPT-4 with temperature $\tau_{\mathrm{ask}}=0.5$. For the search engine, we use Google Scholar, where we collect the top-5 papers that are accessible. The implementation details are explained in an experiment in the appendix. We access papers through publicly available APIs, such as Arxiv, PMC, OpenAccess, PubMed, and our own local database of papers. Lastly, we set the number of sources to consider at each round of gather_evidence to 20, and the count of evidence context to include in the final answer within answer_question to 8.
@@ 6 Limitations
While recent work has been conducted on prompt optimization [75, 76], the complex setting of multiple agents with individual prompts is unsolved. Specifically, the task becomes a non-trivial, bi-level optimization problem. Consequently, discerning the impact of manual prompt adjustments becomes difficult. Thus, it is unlikely our prompts are optimal and it is difficult to assess which pieces of the prompts are necessary.
@@ Acknowledgments
Dual-use (applying technology for both beneficial and detrimental purposes) is an ongoing concern as more powerful models are developed Urbina et al. [77]. We performed rigorous red-teaming questions and did not observe risks for harm that are significantly elevated compared to direct use of the language models used for summarization. Similarly, the LitQA dataset and responses pose no risk for harm.
@@ Abstract:
Provide num_keywords unique keyword searches (one search per line) and year ranges that will find papers to help answer the question. Do not use boolean operators. Make sure not to repeat searches without changing the keywords or year ranges. Make some searches broad and some narrow. The current year is get_year().
@@
We thus evaluate recall, i.e. finding the original paper. Figure 2 shows the cumulative recall curve, where Google Scholar and Semantic Scholar show outstanding ability to retrieve the original paper. Although CORE API [78] does not perform on par with the other two, we include it here as their main contribution to the field of scientific literature is their standardisation of open-access article repositories.
@@ Appendix G Impact of Parametric Knowledge
Write an answer (about 200 words, but can be longer if necessary) for the question below based on the provided context. If the context provides insufficient information and the question cannot be directly answered, reply ‘‘I cannot answer’’. For each part of your answer, indicate which sources most support it via valid citation markers at the end of sentences, like (Example2012).
@@ G.1 Contradicting Information
In the first example, we ask: “Are COVID-19 vaccines effective?”. The goal of this experiment is to evaluate the inclusion of background information that contradicts gathered context. Thus, while the provided context supports the efficacy of covid vaccines, we provide the model with the following background information: “COVID-19 vaccines are known to be ineffective.”
With the contradicting background information, PaperQA responds that it cannot answer. However, if the extra contradicting information is excluded, the model provides an answer:
@@ G.2 Absence of Key Information
As a second example, we ask PaperQA: “What temperature does water turn to plasma?” and provide it with context that is relevant but insufficient to answer the question. Then, we provide the model with the following background information:
“Water turns into plasma under extreme conditions of temperature and pressure. Plasma is the fourth state of matter, beyond solid, liquid, and gas. It’s a super-heated state where atoms are ionized, meaning electrons are stripped from atoms. This typically occurs at temperatures exceeding 20,000 Kelvin.”